1. Determine the center and radius given the equation of a circle in standard form.



1. (x-5)+(y + 4)² = 64


2. (x + 9) +(y-7)² = 121


3. x² + (y+6)² = 4


4. (x-1)² + y² = 1



• II. Convert the equation of a circle in standard form to general form or vice versa.



1. (x-2)² + (y+6)² = 4 2. (x+6)² + (y + 4)² = 36



3. x² + y² + 4x-2y-4=0

1
Expert's answer
2022-03-08T10:39:02-0500

Solution (1)


"\begin{array}{l}
\left( {x - 5} \right) + \left( {y{\rm{ }} + {\rm{ }}4} \right){\rm{ }} = {\rm{ }}64\\
{\left( {x - 5} \right)^2} + {\left( {y - \left( { - 4} \right)} \right)^2} = {8^2}
\end{array}\" Center is "(5,-4)" and Radius "8"



"\begin{array}{l}
{\left( {x + 9} \right)^2}{\rm{ }} + \left( {y - 7} \right){\rm{ }} = {\rm{ }}121\\
{\left( {x - \left( { - 9} \right)} \right)^2}{\rm{ }} + \left( {y - 7} \right){\rm{ }} = {11^2}
\end{array}\" Center is "(-9,7)" and Radius "11"



"\begin{array}{l}
{\left( {x - 0} \right)^2} + {\left( {y + 6} \right)^2} = 4\\
{\left( {x - 0} \right)^2} + {\left( {y - \left( { - 6} \right)} \right)^2} = {2^2}
\end{array}\" Center is "(0,-6)" and Radius "2"



"\begin{array}{l}
{\left( {x - 1} \right)^2} + {\rm{ }}{y^2}{\rm{ }} = 1\\
{\left( {x - 1} \right)^2} + {\left( {y - 0} \right)^2} = {1^2}
\end{array}\" Center is "(1, 0)" and Radius "1"




Solution (2)


"\begin{array}{l}
{\left( {x - 2} \right)^2} + {\rm{ }}{\left( {y + 6} \right)^2} = {\rm{ }}4\\
{x^2} - 4x + 4 + {y^2} + 12y + 36 = 4\\
{x^2} + {y^2} - 4x + 12y + 28 = 0
\end{array}\" The general form is "{x^2} + {y^2} - 4x + 12y + 28 = 0"




"\begin{array}{l}
{\left( {x + 6} \right)^2} + {\left( {y + 4} \right)^2} = 36\\
{x^2} + 12x + 36 + {y^2} + 8y + 16 = 36\\
{x^2} + {y^2} + 12x + 8y + 16 = 0
\end{array}\" The general form is "{x^2} + {y^2} + 12x + 8y + 16 = 0"




"{x^2} + {y^2} + 4x - 2y - 4 = 0" This is already in the general form